P=2L+2W P=23,AND L=7 Solve for W. For some reason Valencia Comm. College thinks that the answer is W=2, I disagree. My answer is 4.5 again and again.

Answer by checkley75(3666)

P=2L+2W
23=2*7+2W
23=14+2W
2W=23-14
2W=9
W=9/2
W=4.5
I'LL HAVE TO AGREE WITH YOU. ASSUMING YOU HAVE THE RIGHT EQUATION & FACTS.